OpenStudy (ny,ny):

Solve the systems of equations using elimination. 4x-y=105 x+7y=-10

OpenStudy (eamier):

if you want to eliminate, make one variable same coefficient. lets take x. so second eq. should multiply by 4. then first eq. minus second eq..

OpenStudy (eamier):

4x+28=-40 second eq should be like this.

OpenStudy (ny,ny):

4(4x-y=105) no I'm sure you have to add the equations, not subtract.

OpenStudy (eamier):

oh its ok to add or subtract in elimination

OpenStudy (eamier):

if you want to add variable y should have same coefficient

OpenStudy (eamier):

we can try both.

OpenStudy (ny,ny):

ok. 4x-y=105 -4(x+7y=-10) -4x-28y=40 right?

OpenStudy (eamier):

this also correct.

OpenStudy (eamier):

you know next step right? just add both of them

OpenStudy (ny,ny):

yeah. but that will leave me with -y-28y=145. what is -y-28y?

OpenStudy (eamier):

-y-28y=-29y.

OpenStudy (ny,ny):

rightttt. thanks. y=-5. right?

OpenStudy (eamier):

yes

OpenStudy (eamier):

sure you know x too

OpenStudy (ny,ny):

yeah i know how to solve for x. I'll give you the answer, tell me if it's right.

OpenStudy (ny,ny):

x=25?

OpenStudy (eamier):

yes. good job

OpenStudy (ny,ny):

thank you!